OpenStudy (anonymous):

Write each function in intercept form. f(x)=x^2+6x+5 f(x)=x^2-2x-8

OpenStudy (aum):

Factor each quadratic expresion.

OpenStudy (anonymous):

Im not sure how to do that.

OpenStudy (aum):

x^2 + 6x + 5 = (x + ?)(x + ?) You need to fill the question marks with numbers such that their sum is the middle coeeficient 6 and their product is the last term 5.

OpenStudy (anonymous):

(x+2)(x+3)?

OpenStudy (aum):

2 amd 3 add up to 5 and multiply to 6. We need ttwo numbers that ddd to 6 and multiply to 5.

OpenStudy (anonymous):

5 and 1

OpenStudy (aum):

yes.

OpenStudy (anonymous):

how do I make it intercept form?

OpenStudy (aum):

That is it. f(x) = x^2 + 6x + 5 = (x + 1)(x + 5)

OpenStudy (anonymous):

oh. Thank you.

OpenStudy (anonymous):

Is it the same way with negatives?

OpenStudy (aum):

Yes find two numbers that add to -2 and multiply to -8. If two numbers multiply to a negative number then one number has to be positive and the other number has to be negative. What are the various factors of 8? List them and you can find the two numbers.

OpenStudy (anonymous):

-4 and 2 are the two numbers

OpenStudy (aum):

correct.

OpenStudy (anonymous):

How did you get -5 and -1 for the first one

OpenStudy (anonymous):

Im not sure if there is another step

OpenStudy (anonymous):

or +5 and +1

OpenStudy (anonymous):

not negative

OpenStudy (anonymous):

I wrote them down positive

OpenStudy (anonymous):

i rwad x intercepts

OpenStudy (aum):

Does the question ask for the x-intercepts? What you have posted just asks for the intercept form of x^2 + 6x + 5.

OpenStudy (anonymous):

oh I understand now

OpenStudy (anonymous):

I got the wrong answer originally

OpenStudy (aum):

pkay.

OpenStudy (anonymous):

How would I do it if it was like f(x)=2x^2-8x+6